OpenStudy (anonymous):

determine the defining equations of f^-1 by interchanging the variables f(x)=x2+2

OpenStudy (anonymous):

y=x^2+2 x=y^2+2 y^2=x-2 y=+/- sqrt(x-2) f^-1=+/- sqrt(x-2)

OpenStudy (anonymous):

thank you very much!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

if i were to plug this in a graphic calculator would i get an inverse function?

OpenStudy (anonymous):

or is it not a function at all unless i make a restriction?

OpenStudy (anonymous):

its not a function unless you make a restriction because the original equation: y=x^2+2 is not a one-to-one function

OpenStudy (anonymous):

ok thank you